OpenStudy (anonymous):

factor 6x^2+4x then solve using the zero-product property.

Directrix (directrix):

6x^2+4x = 0 2x( 3x + 2) = 0 If the product of two factors is 0, then one or both of the factors = 0. 2x = 0 or 3x + 2 = 0 x = 0 or x = -2/3